A graduating student keeps applying for jobs until she has three offers. The




probability of getting an offer at any trial is 0.48.




a. What is the expected number of applications? What is the variance?

1
Expert's answer
2021-12-19T13:22:16-0500

It is a negative binomial distribution: "X\sim NB(s, p)."

Given "s=3, p=0.48."

a)


"E(X)=s\/p"

"=3\/0.48=6.25"

"Var(X)=s(1-p)\/p^2"

"=3(1-0.48)\/(0.48)^2=6.770833"